OpenStudy (waheguru):

9a-2 3a-1 ---- - ------ 4 2

OpenStudy (waheguru):

help me plz

OpenStudy (anonymous):

=0?

OpenStudy (waheguru):

simplaify it

OpenStudy (anonymous):

ah

OpenStudy (waheguru):

plz i really need help

OpenStudy (anonymous):

2(9a-2)-4(3a-1) / 8 = 18a - 4 - 12a + 4 / 8 =6a / 8 =3/4 * a

OpenStudy (waheguru):

how?

OpenStudy (anonymous):

cos i'm alpha and you are beta

OpenStudy (anonymous):

hmm...first i think tat u get the lcm and first simplify the fractions to make it easier...

OpenStudy (waheguru):

ca n u hlep me plz wat r u getting 4 ur answer

OpenStudy (anonymous):

i think its incorrect..but i got 3a-4/4....

OpenStudy (waheguru):

yeaaaa tats what i ot thx

OpenStudy (anonymous):

really? kay...i hope i helped...!